Please solve this problem.
"Angel is twice as old as Gavin.In three years,Angel will be 13 years less than twice as old as Gavin.Find their ages now."

Answer by ankor@dixie-net.com(22740)

Angel is twice as old as Gavin.
a = 2g
:
In three years,Angel will be 13 years less than twice as old as Gavin.
a + 3 = 2(g+3) - 13
a + 3 = 2g + 6 - 13
a = 2g - 7 - 3
a = 2g - 10
:
Find their ages now.
replace a with 2g (from the 1st statement)
2g = 2g - 10
obviously, this is not possible, no solution